OpenStudy (anonymous):

Which of the following algebraic expressions represents the square of a difference?

OpenStudy (goldphenoix):

What's the following?!??! o_O

OpenStudy (anonymous):

(n - 2) (n - 2)2 (n + 2)(n - 2) (n - 2)(n + 2)

OpenStudy (tkhunny):

(n-2) - A difference (n-2)^2 -- A difference, squared -- A square of a difference (n+2)(n-2) = n^2 - 4 -- A difference of squares (n-2)(n+2) = n^2 - 4 -- A difference of squares